# Coverage Axis++: Efficient Inner Point Selection for 3D Shape Skeletonization

### v5

@@ 4. Method
In this section, we present our algorithm for the selection of skeletal points, which involves a set of candidate points denoted as $P=\{p_{i}\}$, and a group of sampled points represented as $S=\{s_{j}\}$ on the surface. Each point is defined by its 3D Cartesian coordinates in $\mathbb{R}^{3}$. In this paper, all inner point candidates are randomly generated inside the volume; See an example in Fig. 2 (b). The following steps outline our algorithm for skeletal point selection.
4.1. Preprocessing
Following (Dou et al., 2022), the first step is to compute the radii $R=\{r_{i}\}$ for the candidate points, where $r_{i}$ is defined as the closest distance of $p_{i}$ to the sampled surface set $S$, i.e.,
(3) $r_{i}:=dist(p_{i},S)=\min_{s_{j}\in S}\|p_{i}-s_{j}\|.$
The set $R$ together with the center points $P$ define the candidate balls. We slightly dilate all the balls by a small factor to their radii, by offset or scale, namely,
(4) $r_{i}^{\prime}=r_{i}+\delta_{r}\quad or\quad r_{i}^{\prime}=r_{i}(1+\delta_{r}),$
leading to a set of dilated balls $B=\{B_{i}\}$, where $B_{i}$ is the ball centered at $p_{i}$ with radius $r_{i}^{\prime}$. According to Eq. 1, we compute the coverage matrix $\mathbf{D}=(d_{ji})\in\mathbb{R}^{m\times n}$, where $d_{ji}\in\{0,1\}$ takes the value $1$ if the surface point $s_{j}$ is covered by the dilated ball $B_{i}$ and 00 otherwise.
4.2. Coverage Score
@@
To achieve compact skeletonization, we aim to select representative skeletal points that capture the key features of the given 3D object.
If the dilated ball $B_{i}$ of a candidate point $p_{i}$ can cover lots of surface samples from the surface point set $S$, $p_{i}$ should be regarded as representative. This idea matches the one used in (Dou et al., 2022) in the sense that if the minimum number of skeletal points is selected by solving SCP, then the dilated ball of each selected point should cover lots of surface samples.
During the implementation of our algorithm, we denote by $P^{+}$ the set of selected skeletal points and $P^{-}$ the set of unselected ones, such that $P=P^{+}\cup P^{-}$ always holds. We also record the uncovered sampled surface points as $S^{\prime}$, which consists of those uncovered surface samples, i.e.,
(5) $S^{\prime}:~{}\{s_{j}\in S\;\big{|}\;s_{j}\notin B_{i},\;\forall p_{i}\in P^{+}\}.$
@@
Then, we assign a coverage score to each $p_{i}\in P^{-}$, which is determined by the number of samples in $S^{\prime}$ that $B_{i}$ covers. With the help of the coverage matrix $\mathbf{D}$ in an offset manner of Eq. 4, we formally define the coverage score of $p_{i}\in P^{-}$ as follows
(6) $Cove_{i}:=\sum_{j:s_{j}\in S^{\prime}}d_{ji}.$
The introduced coverage score promotes a high coverage rate (See Appendix. A.2) of the overall shape to effectively enforce consistency with the given shape structure while favoring the interior points that dominate a larger area, leading to a compact representation of the input shape in the sense of MAT.
4.3. Uniformity Score
@@
The coverage score intrinsically relates the skeletal points with the model surface, we also need another metric to guarantee the uniformity among the selected skeletal points. In other words, the selected points themselves should be relatively uniformly distributed in the space instead of forming groups. For this purpose, we introduce the uniformity score which aims at punishing those candidate points close to the set of currently selected points $P^{+}$. Specifically, for each $p_{i}\in P^{-}$, its uniformity score is defined as
(7) $Unif_{i}:=dist(p_{i},P^{+})=\min_{p_{i^{\prime}}\in P^{+}}\|p_{i}-p_{i^{\prime}}\|.$
@@
4.4. Final Score
We assign a final score to each candidate in $p_{i}\in P^{-}$ by combining its coverage score and uniformity score. We denote the two score vectors as $\mathbf{Cove}=(Cove_{i})$ and $\mathbf{Unif}=(Unif_{i})$, respectively. Since these two scores have different scales (one is in quantity units, and the other is in length units), we apply standardization to them. Specifically, for each $p_{i}\in P^{-}$, we perform the following transformation to $Cove_{i}$ and $Unif_{i}$
(8) $Cove_{i}\leftarrow\frac{Cove_{i}-mean(\mathbf{Cove})}{std(\mathbf{Cove})}$
@@ 4.4. Final Score
where $mean(\cdot)$ and $std(\cdot)$ represent the sample mean and sample standard deviation respectively. Then we add $Unif_{i}$ scaled by a weight parameter $\omega$ to $Cove_{i}$ to derive the final score for $p_{i}$, i.e.,
(9) $Score_{i}:=Cove_{i}+\omega\cdot Unif_{i}.$
After computing $Score_{i}$ for all $p_{i}\in P^{-}$, we select the candidate that achieves the highest score. An overview of the heuristic algorithm is summarized in Algorithm 1.
Input: sampled surface points $S$, candidate skeletal points $P$, dilation factor $\delta_{r}$, weight parameter $\omega$, target number of selected points $|V|$.
@@
      Compute the mean and standard deviation for the two vectors $\mathbf{Cove}=(Cove_{i})$ and $\mathbf{Unif}=(Unif_{i})$;
@@
             Standardize $Cove_{i}$ and $Unif_{i}$ according to Eq. 8;
@@
      Select $p_{i_{k}}\leftarrow\arg\min_{p_{i}\in P^{-}}\{Score_{i}\}$;
@@
4.5. Connectivity of Skeletal Points
After selecting the skeletal points, we establish connections among them to produce the skeletonization result.
Mesh Input
Following (Amenta et al., 2001; Wang et al., 2022b; Dou et al., 2022), we initialize the candidate based on Voronoi diagram VD. After selecting the skeletal points, we can obtain selected points embedded in the Voronoi diagram by nearest projection. We simplify VD using the edge collapse strategy of Q-MAT (Li et al., 2015) following (Dou et al., 2022). The simplification result $\textbf{VD}^{\prime}$ serves as the final MAT of the given mesh input. Note topology preservation (Dey et al., 1998) and mesh inversion avoidance algorithm  (Garland and Heckbert, 1997) are applied during the edge collapse.
Point Cloud Input
As it is hard to preserve the topology during connection establishment for point clouds, we compute an approximation of the MAT connection following (Amenta et al., 2001). We compute the Power diagram PD on the set of $\{B^{\prime}(p_{i},r_{i}^{\prime}),p_{i}\in P^{+}\}\cup\{B(s_{j},\delta_{r}),s_{j}\in S\}$ where $P^{+}$ and $S$ are sets of selected points and surface points, respectively. Then, the connectivity among skeletal points is obtained by extracting edges among the selected inner points on the dual of PD, regular triangulation RT (in other words, weighted alpha complexes). Note that $r^{\prime}$ can be adjusted accordingly during the computation of RT to achieve flexible control over the connection (Dou et al., 2022; Amenta et al., 2001). An overview of our pipeline is shown in Fig. 2.
4.6. Discussion
Complexity
@@ 4.6. Discussion
Algorithm 1 has a polynomial complexity. Specifically, computing $Cove_{i}$ and $Unif_{i}$ costs $O(|S|)$ and $O(|V|)$ operations. Hence summing up for all $p_{i}\in P^{-}$ gives complexity $O(|P|\cdot|S|)$, considering $|V|$ is usually very small compared to $|S|$ and $|P|$. Standardizing $\mathbf{Cove}$ and $\mathbf{Unif}$, as well as sorting, each take $O(|P|)$ operations. Taking into account that the procedure repeats for a maximum of $|V|$ times, the overall complexity of Algorithm 1 can be expressed as $O(|S|\cdot|P|\cdot|V|)$. Fig. LABEL:fig:time_ablation validates this complexity empirically by illustrating the impact of different set sizes $|P|$, $|S|$ and $|V|$ on the running time of Coverage Axis++. In contrast, Coverage Axis requires an exponential cost in the worst case. Although modern solvers accelerate the solving process, long computation time is incurred in the worst case; See Tab. 1.
@@
Table 1. Quantitative comparison on run time and shape approximation error between Coverage Axis and Coverage Axis++. The inputs are meshes and point clouds. Time: Runtime measured in seconds. $|V|$: The number of skeletal points. $\overrightarrow{\epsilon}$: One-sided HD from surface to reconstruction. $\overleftarrow{\epsilon}$: One-sided HD from reconstruction to surface. $\overleftrightarrow{\epsilon}$: Two-sided HD between original surface and reconstruction.

### v7

@@ 4 Method
Our goal is to select skeletal points from a set of candidate points denoted as $P=\{p_{i}\}$, by utilizing a group of sampled points represented as $S=\{s_{j}\}$ on the surface, where each point is defined by its 3D Cartesian coordinates in $\mathbb{R}^{3}$. We present an effective and efficient algorithm for skeletal point selection. We denote by $P^{+}$ the set of selected skeletal points and $P^{-}$ the set of unselected ones, such that $P=P^{+}\cup P^{-}$ always holds. We gradually select a point from $P^{-}$ and add it to $P^{+}$ until the selected points reach the number specified by the user. In this paper, all inner point candidates are randomly generated inside the volume; See an example in Fig. 1 (b). The algorithm for shape skeletonization is outlined through the following steps: scoring each candidate point for point selection and establishing connectivity among skeletal points.
4.1 Coverage Score
@@
To achieve compact skeletonization, we aim to select representative points that capture the key features of the given 3D object by measuring its shape coverage [DLX22] capability.
We first compute the medial radii $R=\{r_{i}\}$ for the candidate points, where $r_{i}$ is defined as the closest distance of $p_{i}$ to the sampled surface set $S$, i.e.,
$r_{i}:=dist(p_{i},S)=\min_{s_{j}\in S}\|p_{i}-s_{j}\|.$ (1)
The set $R$ together with the center points $P$ define the candidate balls. We slightly dilate all the balls by a small factor to their radii by offset, namely,
$r_{i}^{\prime}=r_{i}+\delta_{r},$ (2)
leading to a set of dilated balls $B=\{B_{i}\}$, where $B_{i}$ is the ball centered at $p_{i}$ with radius $r_{i}^{\prime}$. A matrix $\mathbf{D}\in\{0,1\}^{m\times n}$ is introduced by [DLX22], where $m$ and $n$ are total numbers of sampled surface points and candidate skeletal points, respectively. Each entry $d_{ji}\in\{0,1\}$ of $\mathbf{D}$ indicates if the surface point $s_{j}$ is covered by the dilated ball $(p_{i},r_{i}^{\prime})$:
$d_{ji}=\left\{\begin{aligned} 1,&\ \text{if}\ \left\|p_{i}-s_{j}\right\|_{2}\leq r_{i}^{\prime},j=1,...,m,\ i=1,...,n\\ 0,&\ \text{if}\ \left\|p_{i}-s_{j}\right\|_{2}>r_{i}^{\prime},j=1,...,m,\ i=1,...,n.\\ \end{aligned}\right.$ (3)
According to Eq. 3, we have the coverage matrix $\mathbf{D}=(d_{ji})\in\mathbb{R}^{m\times n}$, where $d_{ji}\in\{0,1\}$ takes the value $1$ if the surface point $s_{j}$ is covered by the dilated ball $B_{i}$ and 00 otherwise.
In this paper, we define $S^{\prime}$ as the uncovered surface points, i.e.,
$S^{\prime}:~{}\{s_{j}\in S\;\big{|}\;s_{j}\notin B_{i},\;\forall p_{i}\in P^{+}\}.$ (4)
@@
Then we present the coverage score to each $p_{i}\in P^{-}$, which is determined by the number of samples in uncovered surface points $S^{\prime}$ that $B_{i}$ covers
$Cov_{i}:=\sum_{j:s_{j}\in S^{\prime}}d_{ji}.$ (5)
If the dilated ball $B_{i}$ of a candidate point $p_{i}$ can cover lots of surface samples from the surface point set $S^{\prime}$, $p_{i}$ should be regarded as a representative candidate, which indicates that it has high (local) shape approximation capability. The coverage score favors the medial balls that dominate a larger local volume of the given shape during the selection and thus promotes a high coverage rate (See Appendix. C) of the overall shape to effectively enforce consistency with the input shape.
4.2 Uniformity Score
@@
In addition to the coverage score that intrinsically relates the skeletal points with the model surface, one needs to guarantee uniformity among the selected skeletal points. In other words, the selected points themselves should be relatively uniformly distributed in the space instead of clustering together to avoid redundancy. For this purpose, we introduce the uniformity score, which aims at punishing those candidate points close to the set of currently selected points $P^{+}$. Specifically, for each $p_{i}\in P^{-}$, its uniformity score is defined as
$Unif_{i}:=dist(p_{i},P^{+})=\min_{p_{i^{\prime}}\in P^{+}}\|p_{i}-p_{i^{\prime}}\|.$ (6)
@@
4.3 Centrality Score
We further encourage the centrality of the selected points by introducing a centrality score for each candidate skeletal point. The more central a point is in terms of medial ball positioning, the higher its score. It excludes points too close to the surface to avoid their susceptibility to surface irregularities and noise. Specifically, for each $p_{i}\in P^{-}$, we define its centrality score based on its medial radius by
$Centr_{i}:=-\frac{1}{r_{i}},$ (7)
where $r_{i}$ is the radius of $p_{i}$ defined in Eq. 1.
4.4 Final Score
We assign a final score to each candidate $p_{i}\in P^{-}$ by combining its coverage score, uniformity score, and centrality score. We denote the three score vectors as $\mathbf{Cov}=(Cov_{i})$, $\mathbf{Unif}=(Unif_{i})$, and $\mathbf{Centr}=(Centr_{i})$, respectively. Since these scores have different scales, we apply standardization before integrating them. Specifically, for each $p_{i}\in P^{-}$, we perform the following transformation to $Cov_{i}$, $Unif_{i}$ and $Centr_{i}$
$Cov_{i}\leftarrow\frac{Cov_{i}-mean(\mathbf{Cov})}{std(\mathbf{Cov})},$ (8)
@@ 4.4 Final Score
$Centr_{i}\leftarrow\frac{Centr_{i}-mean(\mathbf{Centr})}{std(\mathbf{Centr})},$
where $mean(\cdot)$ and $std(\cdot)$ represent the sample mean and sample standard deviation respectively. Then we integrate the three scores by summing them up, i.e.,
$Score_{i}:=Cov_{i}+Unif_{i}+Centr_{i}.$ (9)
After computing $Score_{i}$ for all $p_{i}\in P^{-}$, we select the candidate that achieves the highest score. An overview of the heuristic algorithm is summarized in Algorithm 1. An ablation study on different scoring schemes is given in Sec. 6.3.
Input: sampled surface points $S$, candidate skeletal points $P$, dilation factor $\delta_{r}$, target number of selected points $|V|$.
@@
      Compute the mean and standard deviation for the two vectors $\mathbf{Cov}=(Cov_{i})$ and $\mathbf{Unif}=(Unif_{i})$;
@@
             Standardize $Cov_{i}$, $Unif_{i}$ and $Centr_{i}$ according to Eq. 8;
@@
      Select $p_{i_{k}}\leftarrow\arg\max_{p_{i}\in P^{-}}\{Score_{i}\}$;
@@
4.5 Candidate and Surface Point Generation
In this paper, for mesh or point cloud inputs, we always randomly generate the set of candidate points $P$ ($|P|$ = 10000) inside the volume using volumetric sampling [^1] https://www.meshlab.net. The number of surface samples $S$ is $1500$ in all the comparisons (except for Appendix. A). For mesh input, we employ Poisson-disk sampling [CCS12] to get the surface sample. For point cloud input, surface samples are obtained by white noise sampling [J21], and fast winding number [BDS18] is used for inside-outside determination.
4.6 Connectivity of Skeletal Points
4.6.1 Mesh Input
For a mesh input, we aim to build up a connection while preserving the topology of the skeleton, for which we employ the edge collapse strategy of Q-MAT [LWS15] which is based on a Voronoi diagram. Following [ACK01, WWX22, DLX22], we initialize the initial Voronoi diagram VD using the surface points. We then embed the selected points in the Voronoi diagram using the nearest projection, following [DLX22]. We simplify VD using the edge collapse strategy [LWS15]. Specifically, during this process, each edge is checked and collapsed using the edge collapse strategy [LWS15] based on the Spherical Quadric Error Metric [LWS15, TGB13]. We use additional conditions during edge collapse: If both endpoints of an edge to be collapsed are not selected points, we directly apply the edge collapse strategy following [LWS15]. If one endpoint of an edge is a selected point and the other is not, during edge collapse, the selected point is preserved. If both endpoints of an edge are selected points, the edge is preserved and skipped. This process continues until all edges have endpoints that belong to the selected points; See Fig. 2. We adopt topology preservation [DEGN99] and mesh inversion avoidance strategy [GH97] during the collapse. The simplification result $\textbf{VD}^{\prime}$ serves as the final MAT of the given mesh input.
[…]
Figure 2: An illustration of the edge collapse process. The surface is in black, with selected skeletal points highlighted in red. The collapsed edges are shown with dashed lines.
Remark. Notably, not all Voronoi vertices inside the given shape can be used to approximate the medial surface. One has to use Poles (a subset of inside Voronoi vertices) for medial surface approximation according to [ACK01]. However, in practice, we find the collapse over the original Voronoi diagram VD yields relatively stable outcomes for a compact medial surface approximation. Furthermore, the connectivity of skeletal points (edges and facets in MAT) is crucial information in the output MAT results. This connectivity is significantly important for downstream applications such as shape analysis and reconstruction, which we discuss in detail in Appendix E.
4.6.2 Point Cloud Input
Topology preservation during connection establishment for point clouds is challenging. The aforementioned edge collapse techniques [LWS15, PWG19] are not applicable as they require a watertight surface. Thus, we approximate the MAT connection using the Power Diagram following [ACK01]. We compute the Power diagram PD on the set of $\{B^{\prime}(p_{i},r_{i}^{\prime}),p_{i}\in P^{+}\}\cup\{B(s_{j},\delta_{r}),s_{j}\in S\}$ where $P^{+}$, $S$ and $r_{i}^{\prime}=r_{i}+\delta_{r}$ (Eq. 2) are sets of selected points, surface points and dilated radius, respectively. Then, the connectivity among skeletal points is obtained by extracting edges among the selected inner points on the dual of PD, regular triangulation RT (in other words, weighted alpha complexes). In the inset figure, the $P^{+}$ points are depicted in red, while the surface samples are shown in black. The medial balls with original radius $r$ and dilated radius $r^{\prime}$ are illustrated in red and yellow, respectively. Connectivity is established by the connections among selected points of the regular triangulation $\mathbf{RT}$, which are also colored as black solid lines. Additionally, the power plane between the two sites is visualized in blue. This approach can also be applied to mesh inputs, as the computation of the Power Diagram does not depend on the underlying surface being watertight. An overview of our pipeline is shown in Fig. 1.
4.7 Discussion
4.7.1 Complexity
@@
Our method (Algorithm 1) has a polynomial complexity. Specifically, computing $Cov_{i}$ and $Centr_{i}$ costs $O(|S|)$ operations, while $Unif_{i}$ requires $O(|V|)$ operations. Hence summing up for all $p_{i}\in P^{-}$ gives complexity $O(|P|\cdot|S|)$, considering that $|V|$ is usually very small compared to $|S|$ and $|P|$. Standardizing $\mathbf{Cov}$, $\mathbf{Unif}$ and $\mathbf{Centr}$, as well as sorting, each take $O(|P|)$ operations. Taking into account that the procedure repeats for a maximum of $|V|$ times, the overall complexity of Algorithm 1 can be expressed as $O(|S|\cdot|P|\cdot|V|)$. Fig. LABEL:fig:time_ablation validates this complexity empirically by illustrating the impact of different set sizes $|P|$, $|S|$ and $|V|$ on the running time of Coverage Axis++. In contrast, Coverage Axis requires an exponential cost in the worst case. Although modern solvers accelerate the solving process, long computation time is incurred in the worst case; See Tab. 1.
@@
Table 1: Quantitative comparison on run time and shape approximation error between Coverage Axis and Coverage Axis++. The inputs are meshes and point clouds. Time: Runtime measured in seconds. $|V|$: The number of skeletal points. We set the number of selected skeletal points of our method the same as the Coverage Axis [DLX22] for a fair comparison.